Project

General

Profile

NDOS Commissioning Meeting 2011/12/08

Attendance: Jon, Denis, Mat, Mark Mathis, Art, Athans, Jan, Susan, Nick, Andrew, Peter,

Agenda:

  1. Beam and Detector Update
  2. Installation Plans for the next week
  3. DAQ Update and Plans for next week
  4. DCS Update and Plans for next week
  5. AOB

Beam and Detector Report

  • Lost a silicone-coated APD. But the other parylene-coated APDs are still operating fine.
  • Susan and Jan have reported that the fraction of noise-only events has increased over time, but after some consideration, this is not considered to be a problem since the underlying APD distributions (both quantity and positions on the detector) have changed over time. So, there is no indication that there are problems with increasing noise in the NDOS.

Installation plans for next week

  • Leon will be here ~all of next week, and will bring 10 silicone-coated APDs and "n" (~6) parylene-coated APDs, along with new mounting hardware
  • Mat and Leon will install these early next week

DAQ Update and Plans for next week

  • 2-Timing chain test is schedule for tomorrow morning
  • Jan: DCM timing issues, are they understood? Were data lost? Jon: no (although there is a viable hypothesis), and probably yes.
  • DSO: does not require rebooting DCMs!
  • DAQ/DCS: no longer need to re-enable cooling after "Configure Hardware" in RC.
  • Peter:
    • upgrade in BNEVB will allow trigger search, needs a stress test (high trigger rate). Will schedule
    • test of handling very long trigger windows
    • BNEVBs have been improved to be able to handle a disappearing DCM
    • new TDUControl to help catch sync problems
    • total DDS message rate tests

DCS Update and Plans for next week

  • PVArchiver failed around 11/15. Failed SQL writes have been logged to a flat text file, Jon will attempt to add these data into the dB soon, and will send out instructions/documentation to the DCS group on how to recover from such a problem.

AOB

  • Art:
    • some confusion by swing shifters as to whether or not to do a DSO scan. Jon: there should have been no confusion, clear instructions were sent to the shifters. Apparently they did not read those instructions.
    • Doug had to do 5 sequences of start/stop DAQ. Normal? Jon: no.